# xBound: Join Size Lower Bounds

### v1

@@ 8. Conclusion
In this work, we set out to address an issue that was identified more than ten years ago by Leis et al. (Leis et al., 2015): Systems tend to underestimate. And they still do so today. In these ten years, the theory community has brought us exciting new results on join size upper bounds. From the perspective of a real system, replacing one’s own estimator with a new one requires considerable courage, particularly when these estimators still exhibit significant one-sided Q-errors. For this reason, upper bounds are best suited for correcting the original optimizer’s overestimates.
Observing this need, we introduced xBound, the very first framework for provable join size lower bounds. While limited to multi-way joins on the same join key, xBound’s lower bounds are practical enough to both unearth bound violations in modern optimizers, as well as improve, for the first time, their one-sided Q-errors (by capping the estimate with the corresponding lower bound). This could be seen even on a production workload, on which xBound fixed 36.1% of Fabric Data Warehouse’s underestimates. Given these observations, we outlined several open points that are worth exploring from both a research and systems perspective.

### v2

@@ 8. Conclusion
Motivated by the pronounced long tail of CPU resource underestimates observed in production, we introduced in this work xBound, a framework for computing cardinality lower bounds. Although it currently supports only inner (and outer) joins on single keys, preliminary results show that it can correct underestimates and fix end-to-end query execution slowdown in the most painful cases on the challenging open benchmark StackOverflow-CEB in our system, Fabric DW.
We outlined several avenues for generalization, including extending the range of supported query types as well as advancing the underlying theory to further tighten the lower bounds. We hope to spark interest in this problem among both practitioners and theorists.
@@
Acknowledgments. We thank our colleagues Antonio Fernández-Rodríguez, José Alberto López Magaña, and Tej Trivedi for their vital hands-on support for the Fabric Warehouse setup.